X^2 +4y^2=16 2x^2-y^2=12

Hey everyone,


I am struggling to get to the answer for this question. After some rearranging I have managed to rearrange it into a different format but I'm stuck on how to get it into the quadratic formula. 


Any help would be appreciated.


X2+4y2=16

2x2-y2=12 


I got as far as 2(16-4y2)-y2=12
Answer #1

x2+4y2 = 16 ---->1

 2x2-y2=12 ; y2=2x2-12 ----->2

Substituting value of y2 in 1 from 2:

x2+4(2x2-12)=16

9x2-48=16

9x- 54 = 0

9(x2-6)=0

Therefore, x= +61/2 and x= -61/2 satisfy the two equations.


answered by: Saanvi Singh